If in triangle ABC ,a=2b and |A-B|=π/3, Then find angle C

Grade:11

1 Answers

Suraj Kumar
6 years ago
Using formula Tan (A-B)/2=(a-b)/(a+b)×cotc/2We have Tan (pi/3)=(b/3b)×cot C/2€root over((3)=1/3×cotC/2CotC/2=root over 3CotC/2=cot(pi/6)C/2=(pi/6)C=(pi/3)OR C=60°THANKS